OpenStudy (anonymous):

how are questions that are like this solved ? *USING EXPONENT PROPERTIES: simplify the expression. Exponents in your final answer should be positive . 1.(-2st^5u^3)^3

OpenStudy (anonymous):

Power rule: When you raise an exponent to an exponent, you multiply the exponents. For example \[(x^5)^2=x^{10}\]

OpenStudy (anonymous):

can you further explain this ?

OpenStudy (anonymous):

Damn are the equations working here? I don't see what I wrote after for example...

OpenStudy (anonymous):

(x^5)^2=x^(5*2)

OpenStudy (anonymous):

\[-2st^5u^3)^3 = (-2)^3s^3t^{5*3}u^{3*3}\]

OpenStudy (anonymous):

what do those symbols mean?

OpenStudy (anonymous):

Did they show up right yet?

OpenStudy (anonymous):

idk what the symbols in between the exponents mean though?

OpenStudy (anonymous):

For some reason, the site isn't displaying equations correctly, ignore those symbols except for the multiplication signs (*)